5.

Let u = (1, 2, 3) and v = (-2, 1, 2) Find u- v and v u | Need Help? Lm m 1 Talk to a Tutor Read It Show My Work (Requiredy What steps or reasoning did you use? Your work c You can submit show my work an unlimited numbe

Please help me solve the following Linear Algebra problem. Please show work.
Answer #1

u = ( 1 , 2 , 3 ) , v = ( - 2 , 1 , 2 )

u - v = ( 1+ 2 , 2 -1 , 3 - 2 )

= ( 3 , 1 , 1 )

v - u = ( - 2 -1 , 1 - 2 , 2 - 3 )

= ( - 3 , -1 , -1 )